# Optimistic Exploration for Risk-Averse Constrained Reinforcement Learning

### v1

@@ 1 Introduction
In safety critical domains, it is essential for Reinforcement Learning (RL) systems to adhere the safety constraints of the environment [13, 15]. Constrained RL (CRL) has emerged as a framework to address this, to train policies that maximise a reward function whilst satisfying safety constraints placed on a set of cost functions [2, 27]. Uncertainty that arises in many real-world scenarios, however, can pose significant challenges to the safety of CRL policies [11], as the standard CRL approach is to satisfy constraints in expectation. This leaves CRL policies susceptible to constraint violations caused by the inherent randomness of the environment. Because of this, Risk-averse CRL (RaCRL) has gained attention recently as it aims to minimise the probability of rare but catastrophic constraint violations due to environmental randomness [33, 21].
In general, however, introducing risk-aversion leads to a more pronounced trade-off between reward and cost, as it induces conservatism in the policy by reducing the set of feasible constraint satisfying policies [34]. This set of feasible policies will naturally depend on the desired level of risk-aversion but it will also depend on the extent of inherent uncertainty within the environment. Yang et al. [33] propose an off-policy actor-critic RaCRL algorithm called Worst-Case Soft Actor Critic (WCSAC), that makes use of Distributional RL [10, 4] to learn a distributional representation of cost function returns to capture the environment’s inherent uncertainty. Using this representation, WCSAC calculates the Conditional Value at Risk [26] of the policy’s cost function returns (i.e., the worst-case cost returns), to which it aims to maintain below the safety thresholds.
@@
In particular, ORAC constructs an exploratory policy such that at each exploration step it maximises a local upper confidence bound of the state-action reward value function whilst minimising a local lower confidence bound of the risk-averse state-action cost value function. We also develop a weighting scheme for the cost value that increases or decreases at each step should it exceed or fall below the safety constraint value. This way, by leveraging epistemic uncertainty, ORAC bias’ the policy’s actions towards those that may lead to higher reward whilst satisfying costs. ORAC encourages the policy to explore uncertain regions of the environment to discover high reward states whilst still satisfying the safety constraint
@@ 2 Related Work
Specification of safety in RL is a persistently challenging problem [13, 11]. Constrained RL (CRL) has emerged as an effective framework for maximising a reward objective whilst ensuring safety constraints over a separate set of safety objectives are satisfied [2, 25, 27, 22, 35]. Standard CRL is in general, however, neutral to infrequent but catastrophic constraint violations that may arise due to inherent uncertainty within the environment, as safety constraints are satisfied in expectation. Yang et al. [33] aim address this by mitigating risk in CRL and propose an off-policy CRL approach extending Soft Actor-Critic aiming to satisfy CVaR safety thresholds. The authors propose the use of a distributional value function for the safety critic, specifically Implicit Quantile Network architecture [9], allowing for the learning of a expressive representation of the quantile distribution of cost-returns. Kim et al. [20] proposed a trust-region based algorithm that aims to satisfy mean/standard deviation constraints that reduces the uncertainty in constraint satisfaction, but not specifically risk, as mean/standard deviation is not a coherent risk measure. Most recently, Yang et al. [34] proposed an on-policy approach using backward value functions, for non-stationary policies, showing strong constraint satisfaction in multiple environments but at a severe reward-cost trade-off. Recently, Kim et al. [21] propose a general approach to linearise spectral risk-measures, such as CVaR, providing strong computational benefits over the non-linear approximation used in [33].
@@
Exploration is a widely studied area in standard RL, with a diverse range of approaches developed but less so in Constrained RL. Bharadhwaj et al. [5] present a conservative exploration approach that resamples actions from a policy if the conservative expected value estimate of the action exceeds the safety threshold. Wachi et al. [31] present a generalised formulation for safe exploration in constrained RL where the action the agent takes in the is interrupted if it is going to violate constraints, and the environment is reset. This approach is akin to shielding [7], in which the agent is prevented from taking an action that leads to constraint violations. These methods however rely on strict assumptions such as prior knowledge of the underlying MDP, a given safe policy or set of actions or interruptible actions, or may lead to over conservativism in exploration, particularly in complex environments.
@@
We take inspiration from optimistic forms of exploration found in the reward-only RL literature. In the discrete action space, Keramati et al. [19] aim to optimistically explore the environment to quickly learn a risk-averse policy by forming an optimistic return distribution for action selection. In the continuous action setting, Ciosek et al. [8] present an optimistic exploration approach for Actor Critic algorithms such as SAC [16] and TD3 [12] to address conservative exploration that is inherent in these algorithms due to conservative actor updates, inspiring our work. The proposed approach is limited as it only incorporates epistemic uncertainty rather than aleatoric uncertainty which is the primary concern of Risk-Averse RL. Moskovitz et al. [24] present an approach that considers aleatoric uncertainty through learning a distributional reward value function, and switch between being optimistic and conservative in the actor updates of the algorithm, rather than through exploration.
@@ 3.2 Risk Constrained RL
[…]
@@ 3.3 Distributional RL
Using quantile regression, we train the parametrised quantile critic $Z^{\pi}(s,a|\theta)$ by minimising the quantile Huber loss. The Huber loss [17] is defined:
@@ 4.1 Conservative Exploration
Risk-aversion, in general, leads to more conservative exploration of the environment, as it reduces the set of feasible constraint satisfying policies [33, 34]. We present a motivating example of the effect this conservative exploration can have on a risk-averse CRL agent in a GridWorld environment called GuardedMaze [14], shown in Figure 1. In this environment the goal of the agent is to navigate from its initial starting position to the green goal block, choosing one of two paths to reach the goal. The shortest path, passing through the red block, leads to higher reward and in the average case incurs less cost. It is, however, more risky as there is a small probability the cost incurred along this path will be far higher than the safety threshold. Instead, the agent may take the long path, passing through the pink block, and receive less reward but incur a fixed cost that satisfies the safety threshold.
@@ 4.2 Optimistic Exploration
To address the problem of conservative exploration we extend Worst-Case Soft Actor-Critic (WCSAC) [33], by constructing a separate exploratory policy following the principle of Optimism in the Face of Uncertainty (OFU), that explores uncertain parts of the state space to discover high reward states, whilst still satisfying the given safety constraints.
In the case of Soft Actor-Critic algorithms, of which WCSAC is itself an extension, actions are sampled from a parametrised action distribution centred around state-dependent mean values, with state-dependent standard deviations. Inspired by Ciosek et al. [8] we propose a method of shifting this mean value towards action values that maximise a locally optimistic upper bound of the reward value-function and minimise a locally optimistic lower bound of the cost value-function. We begin by defining the learned target policy $\pi_{T}$, which in normal settings is the policy through which the agent explores the environment.
@@
where $\hat{Q}_{R}^{\pi}(s,a)$ is the optimistic upper confidence bound of the reward value function, $\hat{Q}_{C,\alpha}^{\pi}(s,a)$ is the optimistic lower bound of the CVaR of the cost value function, and $\bar{\lambda}$ is the Lagrangian value used to weight the cost value. We define these terms and how each is calculated in detail, in the following sections.
@@ 4.3 Confidence Bounds of Value Functions
Next we define the locally optimistic lower confidence bound of the state-action cost quantile value function, which we denote $\hat{Z}^{\pi}(s,a)$. We extend WCSAC [33] by learning E independent quantile critics $Z^{\pi}(s,a|\theta_{e})$, that share the same network architecture and are each initialised under different sets of weights. During training, each critic learns the cost returns independently and in parallel using the same collected experience.
@@
where $\mu^{k}_{q_{C}}(s,a)=\frac{1}{E}\big{[}\sum_{e=1}^{E}q^{k}_{C,e}(s,a)\big{]}$ is the mean value of the $k$-th quantile values drawn from each critic, and $\sigma_{q_{C}}(s,a)=\sqrt{\frac{1}{E}\sum_{e=1}^{E}\big{(}q^{k}_{C,e}(s,a)-\mu_{q_{C}}(s,a)\big{)}^{2}}$ is the standard deviation of each $k$-th quantile value from the mean quantile value. The locally optimistic lower confidence bound of the state-action cost function $\hat{Z}^{\pi}(s,a)$ is constructed using K optimistic quantiles $q^{\pi}_{c}(s,a)$. Then the locally optimistic CVaR of the cost function can be approximated following [^1] We discuss extensions of this to other approximations of $\hat{Q}^{\pi}_{C,a}(s,a)$ in the supplemental material.:
@@
1:  Initialise: Policy parameters $\phi$, reward critics parameters $\theta_{R_{[1,2]}}$, safety critic parameters $\theta_{C}\}_{i=1}^{E}$, target critics $\bar{\theta}_{R_{[1,2]}}$,$\bar{\theta}_{C}\}_{i=1}^{E}$, risk-aversion level $\alpha$, Lagrangian multiplier $\lambda$, cost threshold $c$, replay buffer $\mathcal{D}$
@@
13:     Update $\phi$, $\theta_{R_{[1,2]}}$, $\theta_{C}\}_{i=1}^{E}$, $\bar{\theta}_{R_{[1,2]}}$,$\bar{\theta}_{C}\}_{i=1}^{E}$ Lagrangian multiplier $\lambda$ following standard WCSAC approach
@@ 5 Experiments
In our experiments we seek to evaluate three main criteria: (a) can conservatism introduced by risk-aversion lead to agents getting stuck in sub-optimal policies? (b) can optimistic exploration introduced by ORAC prevent agents getting stuck in sub-optimal policies? and (c) does ORAC improve reward-cost trade off by finding policies that achieve higher episodic reward, whilst minimising the worst-case cost?
@@ 5.1 Baseline Algorithms
We compare these baseline algorithms against ORAC in three sets of environments, the GuardedMaze Gridworld [14], Safety-Gymnasium [18] and CityLearn [30], across 10 seeds for each agent. We note that in complex environments WCSAC can fail to converge to a risk constraint satisfying policy [21], which we hypothesise is due to the non-linear nature of approximation of risk in the original WCSAC. We therefore extend WCSAC by adopting the approach of Kim et al. [21] to linearise the approximation of risk for Safety Gymnasium and CityLearn environments.  [^2] We provide further details of the practical implementation of both the original implementation of WCSAC and the adoption of the approach of [21], as well as comparisons in Safety Gymnasium tasks, in the supplemental material. We maintain the original implementation of WCSAC in the GuardedMaze environment as applying the approach of [21] failed to find the solution under any conditions, which we believe is due to extra optimisation challenges posed by this approach.
We detail the hyper-parameters used for each algorithms in each environments, in the supplemental material. Specifically we use an upper bound value for the reward value function, $\beta_{R}$, of 3.0 in GuardedMaze and 4.0 in Safety Gymnasium and CityLearn. In GuardedMaze we use a larger optimistic lower bound value for the cost value function, $\beta_{C}$, of 2.0, compared to a smaller bound of 1.0 for all other environments. We use an exploration $\delta$ value of 4.0 that we linearly decay to 0 over the course of training. For fairness of comparison we use the same number of cost value functions for both WCSAC and ORAC, 2 in GuardedMaze and 5 in Safety Gymnasium and CityLearn. We also provide the code used to implement these baselines as well as conduct the experiments, as part of the supplemental material.
@@ 5.2 Optimistic Exploration
Agent Guard Prob. SuccessNo. Steps
@@
PointGoal1 PointButton1 Reward Mean Cost $\text{CVaR}_{0.5}$ $\text{CVaR}_{0.25}$ Reward Mean Cost $\text{CVaR}_{0.5}$ $\text{CVaR}_{0.25}$ SAC Lag 35.1$\pm$0.6 15.5$\pm$5.1 25.4$\pm$6.4 31.0$\pm$6.9 14.6$\pm$4.7 19.7$\pm$4.3 29.4$\pm$6.3 36.1$\pm$8.6 WCSAC-0.5 33.8$\pm$0.5 5.0$\pm$1.1 8.7$\pm$2.0 11.9$\pm$3.2 14.2$\pm$3.5 6.7$\pm$2.4 11.9$\pm$2.8 16.5$\pm$6.0 ORAC-0.5 34.2$\pm$1.2 4.9$\pm$0.9 8.9$\pm$1.5 12.4$\pm$2.0 13.9$\pm$2.7 5.4$\pm$2.1 9.6$\pm$3.5 13.2$\pm$5.2 WCSAC-0.25 32.1$\pm$1.3 3.9$\pm$1.5 7.2$\pm$2.5 10.6$\pm$5.4 10.1$\pm$3.7 4.7$\pm$1.1 8.3$\pm$1.5 11.4$\pm$2.8 ORAC-0.25 32.8$\pm$2.4 3.9$\pm$1.1 7.2$\pm$1.9 10.0$\pm$2.9 11.2$\pm$3.1 4.7$\pm$1.8 8.7$\pm$3.0 12.7$\pm$4.7
@@
We detail the results of these evaluations in Table 1. The risk-neutral agent behaves as expected and converges to the short path under both conditions, for each seed. The risk-averse baseline WCSAC, however, only converges to the risk-averse long path 60% of the time under the 0.15 probability setting and even less so under the 0.1% at just 30%, indicating WCSAC insufficiently explores the environment to find the long path. In contrast, ORAC converges to the risk-averse long path 100% of the time under the 0.15 probability setting and 80% of the time in the 0.1 probability setting, demonstrating clearly the improved exploration of ORAC. We measure the number of steps taking for the algorithms to converge to the long path solution, measuring failed solutions as 500K steps. We note that ORAC leads to faster convergence to the long path under both Guard probability settings, indicating ORAC leads to more sample efficient exploration over the baseline.

### v2

@@ 1 Introduction
In safety critical domains, it is essential for Reinforcement Learning (RL) systems to adhere to the safety constraints of the environment [13, 15]. Constrained RL (CRL) has emerged as a framework to address this, by training policies to maximise a reward function whilst satisfying safety constraints over a set of cost functions [2, 28]. Uncertainty that arises in many real-world scenarios, however, can pose significant challenges to the safety of CRL policies [11], as the standard CRL approach is to satisfy constraints in expectation. This leaves CRL policies susceptible to constraint violations caused by the inherent randomness of the environment. Because of this, Risk-averse CRL (RaCRL) has gained attention recently as it aims to minimise the probability of rare but catastrophic constraint violations due to environmental randomness [34, 21].
In general, however, introducing risk-aversion leads to a more pronounced trade-off between reward and cost, as it induces conservatism in the policy by reducing the set of feasible constraint satisfying policies [35]. This set of feasible policies will naturally depend on the desired level of risk-aversion but it will also depend on the extent of inherent uncertainty within the environment. Yang et al. [34] propose an off-policy actor-critic RaCRL algorithm called Worst-Case Soft Actor Critic (WCSAC), that makes use of Distributional RL [9, 4] to learn a distributional representation of cost function returns to capture the environment’s inherent uncertainty. Using this representation, WCSAC calculates the Conditional Value at Risk (CVaR) [27] of the policy’s worst-case cost returns. Through this, WCSAC learns a policy that, at deployment time, maintains worst-case cost returns below the safety thresholds.
@@
In particular, ORAC constructs an exploratory policy such that at each exploration step it maximises a local upper confidence bound of the state-action reward value function whilst minimising a local lower confidence bound of the risk-averse state-action cost value function. We also develop a weighting scheme for the cost value that increases or decreases at each step should it exceed or fall below the safety constraint value. This way, by leveraging epistemic uncertainty, ORAC bias’ the policy’s actions towards those that may lead to higher reward states whilst minimising worst-case costs. ORAC encourages the policy to explore uncertain regions of the environment to discover high reward states whilst still encouraging satisfaction of the safety constraint.
@@ 2 Related Work
Specification of safety in RL is a persistently challenging problem [13, 11]. Constrained RL (CRL) has emerged as an effective framework for maximising a reward objective whilst ensuring satisfaction of constraints over a set of safety objectives [2, 26, 28, 23, 36]. Standard CRL is in general, however, neutral to infrequent but catastrophic constraint violations that may arise due to inherent uncertainty within the environment, as safety constraints are satisfied in expectation. Yang et al. [34] aim to mitigate risk in CRL and propose an off-policy CRL approach, extending Soft Actor-Critic to CVaR safety thresholds. The authors propose the use of a distributional value function for the safety critic, specifically the Implicit Quantile Network architecture [10], allowing for the learning of a expressive representation of the quantile distribution of cost-returns. Kim et al. [20] proposed a trust-region based algorithm that aims to satisfy mean/standard deviation constraints, reducing uncertainty in constraint satisfaction, but not specifically risk, as mean/standard deviation is not a coherent risk measure. Yang et al. [35] proposed an on-policy approach using backward value functions, for non-stationary policies, showing strong constraint satisfaction in multiple environments but at a severe reward-cost trade-off. Recently, Kim et al. [21] propose a general approach to linearise spectral risk-measures, such as CVaR, providing strong computational benefits over the non-linear approximation used in [34].
@@
Exploration is a widely studied area in standard RL, but is less studied in Constrained RL. Bharadhwaj et al. [5] present a conservative exploration approach that resamples actions from a policy if the conservative expected value estimate of the action exceeds the safety threshold. Wachi et al. [32] present a generalised formulation for safe exploration in constrained RL where the action the agent takes in the environment is interrupted if it is going to violate constraints, and the environment is then reset. This approach is akin to shielding [7], in which the agent is prevented from taking an action that leads to constraint violations. These methods, however, rely on strict assumptions such as prior knowledge of the underlying MDP, access to a safe policy or set of actions, and may lead to overly conservative exploration, particularly in complex environments.
@@
We take inspiration from optimistic exploration strategies found in reward-only RL literature. In the discrete action space, Keramati et al. [19] aim to optimistically explore the environment to quickly learn a risk-averse policy by forming an optimistic return distribution for action selection. In the continuous action setting, Ciosek et al. [8] present an optimistic exploration approach for Actor Critic algorithms such as SAC [16] and TD3 [12] to address conservative exploration that is inherent in these algorithms due to conservative actor updates, inspiring our work. The proposed approach is limited as it only incorporates epistemic uncertainty rather than aleatoric uncertainty which is the primary concern of Risk-Averse RL. Moskovitz et al. [25] present an approach that considers aleatoric uncertainty through learning a distributional reward value function, and switch between being optimistic and conservative in the actor updates of the algorithm, rather than through exploration. Liu et al. [22] propose an optimistic exploration strategy that considers both sources of uncertainty. Treating aleatoric uncertainty as a form of risk, the agent is guided away from regions of high aleatoric uncertainty. The authors approach is different to our work in its aims, namely it is focused on risk-neutral maximisation of reward in reward-only MDP settings, whereas our work aims to maximise reward whilst ensuring risk-based satisfaction of cost constraintsr, balancing optimism with explicit constraint satisfaction.
@@ 3.2 Risk Constrained RL
$\mathcal{R}_{\sigma}(X)\vcentcolon=\int_{0}^{1}F_{X}^{-1}(u)\sigma(u)du$ (4)
where $\sigma$ is an increasing function $\sigma\geq 0,\int_{0}^{1}\sigma(u)du=1$. In the case of CVaR, $\sigma$ is defined as:
@@ 3.3 Distributional RL
Using quantile regression, we train the parametrised quantile critic $Z^{\pi}(s,a|\theta)$ by minimising the quantile Huber loss. The Huber loss [17], is defined:
@@ 4.1 Conservative Exploration
Risk-aversion, in general, leads to more conservative exploration of the environment, as it reduces the set of feasible constraint satisfying policies [34, 35]. We present a motivating example of the effect this conservative exploration can have on a risk-averse CRL agent in a GridWorld environment called GuardedMaze [14], shown in Figure 1. In this environment the agent must navigate from its initial starting position to the green goal block, choosing one of two paths to reach the goal. The shortest path, passing through the red block, leads to higher reward and in the average case incurs less cost. It is, however, more risky as there is a small probability the cost incurred along this path will be far higher than the safety threshold. Instead, the agent may take the long path, passing through the pink block, and receive less reward but incur a fixed cost that satisfies the safety threshold.
@@ 4.2 Optimistic Exploration
To address the problem of conservative exploration we extend Worst-Case Soft Actor-Critic (WCSAC) [34], by constructing a separate exploratory policy following the principle of Optimism in the Face of Uncertainty (OFU), that explores uncertain parts of the state space to discover high reward states, whilst still encouraging satisfaction of the given safety constraints.
In the case of Soft Actor-Critic algorithms, of which WCSAC is itself an extension, actions are sampled from a parametrised action distribution centred around state-dependent mean values, with state-dependent standard deviations. Inspired by Ciosek et al. [8] we propose a method of shifting this mean value towards action values that maximise a locally optimistic upper bound of the reward value-function and minimise a locally optimistic lower bound of the cost value-function. We begin by defining the learned target policy $\pi_{T}$, which in normal settings is the policy through which the agent explores the environment:
@@
fwhere $\hat{Q}_{R}^{\pi}(s,a)$ is the optimistic upper confidence bound of the reward value function, $\hat{Q}_{C,\alpha}^{\pi}(s,a)$ is the optimistic lower bound of the CVaR of the cost value function, and $\bar{\lambda}$ is the Lagrangian value used to weight the cost value. We define these terms and how each is calculated in detail, in the following sections.
@@ 4.3 Confidence Bounds of Value Functions
Next we define the locally optimistic lower confidence bound of the state-action cost quantile value function, denoted $\hat{Z}^{\pi}(s,a)$. We extend WCSAC [34] by learning E independent quantile critics $Z^{\pi}(s,a|\theta_{e})$, that share the same network architecture and are each initialised under different sets of weights. During training, each critic learns the cost returns independently and in parallel using the same collected experience.
@@
where $\mu^{k}_{q_{C}}(s,a)=\frac{1}{E}\big{[}\sum_{e=1}^{E}q^{k}_{C,e}(s,a)\big{]}$ is the mean value of the $k$-th quantile values drawn from each critic, and $\sigma_{q_{C}}(s,a)=\sqrt{\frac{1}{E}\sum_{e=1}^{E}\big{(}q^{k}_{C,e}(s,a)-\mu^{k}_{q_{C}}(s,a)\big{)}^{2}}$ is the standard deviation of each $k$-th quantile value from the mean value of the $k$-th quantile value. The locally optimistic lower confidence bound of the state-action cost function $\hat{Z}^{\pi}(s,a)$ is constructed using K optimistic quantiles $q^{\pi}_{c}(s,a)$. Then the locally optimistic CVaR of the cost function can be approximated following [^1] We discuss extensions of this to other approximations of $\hat{Q}^{\pi}_{C,a}(s,a)$ in the supplemental material.:
@@
1:Initialise: Policy parameters $\phi$, reward critics parameters $\theta_{R_{[1,2]}}$, safety critic parameters $\theta_{C}\}_{i=1}^{E}$, target critics $\bar{\theta}_{R_{[1,2]}}$$\bar{\theta}_{C}\}_{i=1}^{E}$, risk-aversion level $\alpha$, Lagrangian multiplier $\lambda$, cost threshold $c$, replay buffer $\mathcal{D}$
@@
13:  Update $\phi$, $\theta_{R_{[1,2]}}$, $\theta_{C}\}_{i=1}^{E}$, $\bar{\theta}_{R_{[1,2]}}$$\bar{\theta}_{C}\}_{i=1}^{E}$ Lagrangian multiplier $\lambda$ following standard WCSAC approach
@@ 5 Experiments
In our experiments we seek to evaluate three criteria: (a) can conservatism introduced by risk-aversion lead to agents getting stuck in sub-optimal policies? (b) can optimistic exploration introduced by ORAC prevent agents getting stuck in sub-optimal policies? and (c) does ORAC improve reward-cost trade off by finding policies that achieve higher episodic reward, whilst minimising worst-case costs?
@@ 5.1 Baseline Algorithms
We compare these baseline algorithms against ORAC in three sets of environments, the GuardedMaze Gridworld [14], Safety-Gymnasium [18] and CityLearn [31], across 10 seeds for each agent. We note that in complex environments WCSAC can fail to converge to a risk constraint satisfying policy [21], which we hypothesise is due to the non-linear nature of approximation of risk in the original WCSAC. We therefore extend WCSAC by adopting the approach of Kim et al. [21] to linearise the approximation of risk for Safety Gymnasium and CityLearn environments  [^2] We provide further details of the practical implementation of both the original and linear approximation of risk [21] for WCSAC, comparisons of each approach and ablations of the introduced hyperparameters in supplemental material. We maintain the original implementation of WCSAC in the GuardedMaze environment as applying the approach of [21] failed to find the solution under any conditions, which we believe is due to extra optimisation challenges posed by this approach.
We detail the hyper-parameters used for each algorithms in each environments, in the supplemental material. Specifically we use an upper bound value for the reward value function, $\beta_{R}$, of 3.0 in GuardedMaze and 4.0 in Safety Gymnasium and CityLearn. In GuardedMaze we use a larger optimistic lower bound value for the cost value function, $\beta_{C}$, of 2.0, compared to a smaller bound of 1.0 for all other environments. We use an exploration $\delta$ value of 4.0 that we linearly decay to 0 over the course of training. For fairness of comparison we use the same number of cost value functions for both WCSAC and ORAC, 2 in GuardedMaze and 5 in Safety Gymnasium and CityLearn.
@@ 5.2 Optimistic Exploration
Agent Guard Prob. Success% No. Steps
@@
[…]
@@
We detail the results of these evaluations in Table 1. The risk-neutral agent behaves as expected and converges to the short path under both conditions, for each seed. The risk-averse baseline WCSAC, however, only converges to the risk-averse long path 60% of the time under the 0.15 probability setting and even less so under the 0.1% at just 30%, indicating WCSAC insufficiently explores the environment to find the long path. In contrast, ORAC converges to the risk-averse long path 100% of the time under the 0.15 probability setting and 80% of the time in the 0.1 probability setting, demonstrating clearly the improved exploration of ORAC. We measure the number of steps taken for the algorithms to converge to the long path solution, measuring failed solutions as 500K steps. We note that ORAC leads to faster convergence to the long path under both Guard probability settings, indicating ORAC leads to more sample efficient exploration over the baseline.